Q.1 A body is moved along a str. line by a machine delivering a constant power .  The distance travelled by the body in time t is proportional to
  1. t ^3/4
  2. t^3/2
  3. t^1/4
  4. t^1/2                                                   Q2. The potential energy of a 1 kg particle free to move along the x – axis is given by ( X^4/4 – X^2/2) ,then the maximuim speed in m/s is                    a 3/ squareroot 2                                  b. 2^ ½                                             c. 2            

Grade:11

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Dear student,

P = F*V = constt = k
m *a * V = k
m *V2/t = k
m *s2/t3 = k
S = k’ t3/2
….….….….….….….….….….….….….….….….….….….….….….….….….…....
Thanks.